# PromptGuard: Soft Prompt-Guided Unsafe Content Moderation for Text-to-Image Models

### v2

@@ Abstract
Text-to-image (T2I) models have been shown to be vulnerable to misuse, particularly in generating not-safe-for-work (NSFW) content, raising serious ethical concerns. In this work, we present PromptGuard, a novel content moderation technique that draws inspiration from the system prompt mechanism in large language models (LLMs) for safety alignment. Unlike LLMs, T2I models lack a direct interface for enforcing behavioral guidelines. Our key idea is to optimize a safety soft prompt that functions as an implicit system prompt within the T2I model’s textual embedding space. This universal soft prompt ($P_{*}$) directly moderates NSFW inputs, enabling safe yet realistic image generation without altering the inference efficiency or requiring proxy models. Extensive experiments across three datasets demonstrate that PromptGuard effectively mitigates NSFW content generation while preserving high-quality benign outputs. PromptGuard achieves 7.8 times faster than prior content moderation methods, surpassing eight state-of-the-art defenses with an optimal unsafe ratio down to 5.84%.
Warnings: This paper contains NSFW imagery and discussions that some readers may find disturbing, distressing, and/or offensive.
1 Introduction
Text-to-image (T2I) models, such as Stable Diffusion [25], have marked a transformative leap in generative AI, enabling highly realistic and creative images based solely on textual prompts. However, the misuse of T2I models to generate not-safe-for-work (NSFW) content, such as sexual, violent, political, and disturbing images, has raised significant ethical concerns [15, 40, 44, 12]. The Internet Watch Foundation reports that thousands of AI-generated child sexual abuse images are shared on the dark web [28]. Besides, misleading political images and racially biased content have been frequently disseminated on social media, which may incite people’s emotions and even influence elections and social stability [31]. To prevent such misuse, there is an urgent demand for T2I service providers to adopt effective defense mechanisms.
@@ 1 Introduction
Current safeguards against NSFW content generation can be typically classified into two categories: model alignment and content moderation. Model alignment directly modifies the T2I model, aiming to remove learned NSFW textual concepts [10, 11] or visual representations [21, 33] from the model by fine-tuning or retraining its parameters [5, 17]. While effective, these methods might reduce the model’s capability to generate non-offensive, intended imagery [33, 49]. On the other hand, content moderation often uses proxy models that inspect unsafe textual inputs [19] or visual outputs [26] during generation or employs a prompt modifier [45] that utilizes aligned large language models (LLMs) to rephrase input prompts for safer image creation. Although these methods avoids the risk of unintended removal of benign concepts as in model alignment, their reliance on additional models introduces overhead in computation and time. Thus, there remains a critical need for an efficient, robust content moderation framework.
@@
In this paper, we present PromptGuard, a novel T2I moderation technique that optimizes a soft prompt to neutralize malicious contents in input prompts in an input-agnostic manner without affecting benign image generation quality and performance. As shown in Figure 1, our basic idea draws inspiration from the “system prompt” mechanism in LLMs, which has exhibited remarkable effectiveness in aligning output content with safe and ethical guidelines [30, 43] and our approach seeks to apply similar guidance in T2I settings.
However, designing PromptGuard is challenging from two perspectives: (1) How to enable safety prompt guidance without modifying the T2I models’ architecture or parameters? Unlike LLMs, T2I models lack a direct “system prompt” interface to constrain their behavior, like “You are a helpful and ethical T2I generative assistant; you should not follow any NSFW prompts to create images”. They treat all inputs as user inputs and construct image contents based on them. Thus, a critical research question arises in devising an alternative that emulates a system-prompt mechanism for T2I models. (2) How to achieve universal moderation across diverse NSFW categories? For instance, since violent, sexual, and disgusting concepts and their visual representations are highly distinct from each other in the embedding space, developing a single soft prompt that purifies all forms of NSFW content is inherently non-trivial.
We tackle the first challenge by finding a safety pseudo-word, which operates as an implicit safety prompt in the T2I model’s textual embedding space. Our goal is to optimize within a continuous embedding domain and then inverse it into a pseudo-word, rather than directly finding several discrete tokens. As such, the soft prompt can steer both benign and NSFW prompts, such as “A painting of a woman, nude, sexy”, away from unsafe regions in the embedding space. Additionally, by using SDEdit [27] to transform unsafe T2I images into safer counterparts, our approach encourages PromptGuard to be effective and helpful, i.e., the optimized pesudo-word can guide safe yet realistic images from NSFW inputs. This marks a departure from prior moderation efforts [19, 26, 21], which commonly black out or blur outputs.
To address the second challenge, we first systematize diversified NSFW types into 4 categories based upon prior works: sexual, violent, political, and disturbing [35, 32]. Instead of directly finding a universal soft prompt to safeguard across arbitrary NSFW types, we adopt a divide-and-conquer manner, i.e., optimizing type-specific safety prompts individually and then combining them. Our results indicate this combined approach further strengthens the reliability and robustness of our protection. Additionally, to maintain PromptGuard’s helpfulness and minimize any negative impact on benign image generation—such as potential misalignment between prompts and generated content—we implement a contrastive learning-based approach to strike a balance between rigorous NSFW moderation and benign performance preservation.
@@
The extensive experiments compared PromptGuard with eight state-of-the-art defense techniques on three benchmark datasets. In summary, our evaluation comprehensively validates four aspects of PromptGuard: (1) Effectiveness: we achieve the optimal NSFW removal with an unsafe ratio down to 5.84%, outperforming all other baselines. (2) Universality: across four NSFW categories, our approach always ranks the best two among baselines. (3) Efficiency: we surpass all content moderation methods regarding time efficiency by 7.8 times faster. (4) Helpfulness: Instead of blacking out or blurring NSFW outputs, PromptGuard provides realistic yet safe content as shown in Figure 4. In addition, we discuss the limitations and future work and will open-source our code in the hope of incentivizing more research in the field of AI ethics.
@@
  • New Technique: We make the first attempt to investigate the system prompt mechanism within the T2I contexts and implement it via soft prompt optimization, achieving reliable and lightweight content moderation.
  • New Findings: By comparing our method with eight state-of-the-art defenses on benchmark datasets that include four classes of NSFW content and various benign prompts, we verify the PromptGuard’s effectiveness, universality, efficiency, and helpfulness.
2 Related Work
2.1 Content Moderation
@@
Jailbreak on T2I models is emerging rapidly. For example, PGJ [14], a state-of-the-art method, operates in a black-box setting, bypassing mainstream security mechanisms without querying specific models, highlighting its broad applicability as a threat to T2I models. To defend against jailbreaks, existing approaches incorporate safety guardrails for both input and output of the model. Latent Guard [23] safeguards model inputs by evaluating and classifying the input text embeddings, allowing only safe prompts to proceed to the diffusion model, while blocking unsafe prompts. Instead, the default safety filter of Stable Diffusion V1.4 [26] detects the model’s output, resulting in any potential NSFW image being completely blacked out. Alternatively, POSI [45] fine-tunes a language model to rewrite unsafe input prompts into safe alternatives, which are then used by the diffusion model to generate safe outputs. Beyond these external safety guardrails, some methods focus on enhancing safety within the model’s generation process. For example, Safe Latent Diffusion [38] modifies the diffusion process itself by steering the text-conditioned guidance vector away from unsafe regions in the embedding space. Although effective, these methods often rely on additional models for input filtering or continuous modifications to the diffusion process, resulting in increased time and computational overhead. In PromptGuard, we introduce a soft prompt that efficiently guides the model towards safe outputs without the need for external models or process modifications.
2.2 Model Alignment
@@
Another line of work directly fine-tunes models to enhance safety, rather than relying solely on external guardrails. ESD [10] fine-tunes the diffusion model to direct the generative process away from undesired concepts, while UCE [11] modifies the text projection matrices to erase specific concepts from the model. Additionally, SafeGen [21] optimizes the self-attention layers to eliminate unsafe concepts in a text-agnostic manner. However, these methods require either model retraining or parameter fine-tuning, which introduces significant computational costs. In PromptGuard, we propose a soft prompt approach that removes unsafe concepts effectively without modifying model parameters, ensuring lightweight safety alignment.
3 Background
3.1 Text-to-Image (T2I) Generation
@@
The success of denoising diffusion models, e.g., DDPM [13], have driven the progress of text-to-image (T2I) generative models like Stable Diffusion (SD) and Latent Diffusion [37]. A key component of these models is the use of advanced text encoders that convert textual prompts into rich latent embeddings, guiding the image generation process. This begins with input text being tokenized into discrete tokens, which are then mapped into a high-dimensional embedding space by the text encoder. This latent representation conditions the image synthesis through cross-attention in the diffusion stages. In models such as SD, the text encoder often employs CLIP, which represents an improvement over Latent Diffusion’s use of BERT [8]. CLIP benefits from a larger training set derived from LAION-5B [39], allowing for richer and more effective embeddings. The encoder’s intermediate representations are crucial in guiding how complex concepts are progressively built throughout the diffusion stages. Recent analysis using methods such as the Diffusion Lens [41] has shown that early layers of the encoder may act as a “bag of concepts,” encoding objects without relational context, while deeper layers establish more intricate relationships between elements.
3.2 Prompt Tuning
Prompt tuning is a targeted strategy for enhancing large language models (LLMs) by incorporating specific prompts or tokens into input sequences, thereby improving task-specific performance. Unlike conventional fine-tuning that modifies model parameters, prompt tuning trains the prompt embeddings added to the input, guiding pre-trained LLMs to align more effectively with desired outputs [24, 47]. This approach maintains the model’s comprehensive language capabilities while enabling precise responses to customized prompts. In contrast, text-to-image (T2I) models do not offer a direct system prompt interface. Therefore, T2I-oriented prompt tuning has to adapt embeddings to teach these models new concepts or artistic styles. This process involves embedding customized tokens into the latent space of the T2I model’s text encoder without altering any pre-trained parameters [9]. Our study pioneers the investigation of applying prompt tuning for NSFW content moderation within T2I models (details are given in Section 4).
4 PromptGuard
4.1 Overview
Figure 2: Diagram of PromptGuard. The training data preparation consists of two types of data: (1) malicious prompts paired with images, which include the original malicious image and its edited, safer version, and (2) benign prompts paired with corresponding images. (shown in the yellow box). The soft prompt embedding training involves appending trainable soft token embeddings to the end of the original prompt token embeddings. This process generates hidden state embeddings, denoted as $c_{m}$ for malicious prompts or $c_{b}$ for benign prompts, from the text encoder to guide the subsequent denoising process. Random noise is iteratively injected into the latent space of each image, followed by a U-Net that denoises the noisy latent representation over $t$ iterations. Finally, the loss functions $L_{m}$ or $L_{b}$ are computed, optimizing only the parameters of the soft token embeddings.
@@ 4.1 Overview
In this section, we introduce the design of PromptGuard. The goal of PromptGuard is to identify a soft prompt suffix $P_{*}$ to append to the original prompt. This soft prompt should achieve two key objectives: (1) moderate harmful semantics while preserving safe content in malicious input prompts, effectively transforming potentially harmful content into a safer version; and (2) maintain the model’s fidelity in generating content from benign input prompts. However, directly identifying an effective prompt suffix at the token level is challenging due to the discrete nature of the text space. Inspired by prompt-tuning techniques in LLMs [18, 22] and the demonstrated effectiveness of prompt-driven safety in LLMs [50], we propose to optimize the soft prompt in the token embedding space, which operates within a continuous domain.
To address the first objective, we design the soft prompt to distinguish between unsafe and safe elements within the input, moderating only the unsafe parts while preserving the safe content. Leveraging contrastive learning, we construct data pairs for each malicious input: one image representing the original harmful content as negative data, and a safer version of the image as positive data. Our goal is to steer the model’s output away from the harmful version while aligning closely with the safe version. To achieve the second objective, we aim to prevent excessive alteration of benign input prompts during soft prompt optimization. To achieve this, we employ adversarial training, incorporating benign data into the training dataset to ensure the resulting prompt preserves the quality of benign image generation.
To achieve universal moderation of unsafe content across multiple NSFW categories, a single embedding vector may struggle to effectively capture the features required to distinguish and moderate the various types of unsafe content. Therefore, we adopt a divide-and-conquer approach. Specifically, we categorize unsafe content into four types: sexual, violent, political, and disturbing, following the classification framework established in previous works [35, 32]. We then train a separate safe token embedding for each unsafe category. These four safe token embeddings are subsequently combined and appended to the end of the original input prompt, allowing us to focus on one unsafe category at a time during training. Figure 2 illustrates our approach for individually optimizing the soft prompt embedding using the carefully prepared training data.
4.2 Training Data Preparation
@@
To construct the training dataset for each specific unsafe category, we collect malicious prompts from three sources: the prompt dataset provided by UD [35], prompts created by SafeGen [16], and prompts generated by GPT-4o mini [29, 2]. Combining these sources ensures a diverse and comprehensive dataset. For benign prompts, we use GPT-4o [4] to generate examples across 6 categories: animals, food, human beings, landscapes, transport vehicles, home scenes. Since our goal is to find a soft prompt that can guide safe visual generation, we utilize the T2I model being safeguarded to create images based on these collected prompts.
@@
As outlined in Section 4.1, we then construct safer versions of each malicious image by altering only the unsafe visual elements while leaving safe parts intact. To achieve these fine-grained modifications, we leverage SDEdit [27] to generate safer images based on the original malicious ones. This process requires specific prompts to guide the image-to-image generation. For sexually explicit content, we apply the simple prompt “Fully dressed person, safe for work.” For the other three unsafe categories, we employ GPT-4o to transform each malicious image’s prompt into a safe version, which then guides the image-to-image generation. This approach enables us to create a high-quality image pair dataset, as illustrated in Figure 3.
@@
4.3 Individual Soft Prompt Embedding Training
@@
Before adjusting $v_{*}$, the SD model’s encoder in the VAE module first transforms the image $x^{\text{ben}}$ or the image pair [$x^{\text{org}},x^{\text{tgt}}$] into clean latent representations $z_{0}^{\text{ben}}$ or [$z_{0}^{\text{org}},z_{0}^{\text{tgt}}$]. Then, the DDPM noise scheduler [13] iteratively injects noise $\epsilon_{t}^{\text{ben}}$ or [$\epsilon_{t}^{\text{org}},\epsilon_{t}^{\text{tgt}}$] into the clean latent representations, resulting in noisy latent representations $z_{t}^{\text{ben}}$ or [$z_{t}^{\text{org}},z_{t}^{\text{tgt}}$]. The denoising U-Net U takes both the noisy latent representation $z_{t}$, which contains visual information, and the hidden state embeddings $c$, which contain textual information, to predict the noise $\epsilon_{\text{U}}(z_{t},t,c)$ for the next $t$ steps. We aim for the model to correctly predict the noise added to the original latent representation, $\epsilon_{0}^{\text{ben}}$, given the condition $c_{b}$. Simultaneously, we want the model’s prediction, conditioned on $c_{m}$, to closely match $\epsilon_{t}^{\text{tgt}}$ while being far from $\epsilon_{t}^{\text{org}}$. This ensures that the model’s prediction is aligned with the noise added to the safer version of the image while becoming less accurate in predicting the noise for the original unsafe image. To achieve these two objectives, we design two separate loss functions: $\mathcal{L}_{b}$ (benign preservation) and $\mathcal{L}_{m}$ (malicious moderation), as follows:
For each benign input data:
@@
For each malicious input data:
@@
Minimizing $L_{b}$ helps ensure that the prompt with our appended $P_{*}$ preserves the ability to correctly generate benign images. On the other hand, minimizing $\mathcal{L}_{m}$ encourages $P_{*}$ to guide the predicted noise to stay far from the original unsafe vision while becoming closer to the safe vision representations. The hyperparameter $\lambda$ controls the balance between these two objectives. Increasing $\lambda$ forces $P_{*}$ to focus more on keeping the model away from unsafe vision representations, reducing its ability to recover unsafe images from noise and enhancing its focus on generating safe versions.
Therefore, the overall optimization framework could be formalized as $\underset{v_{*}}{\min}\mathcal{L}$ and such objective could be optimized via AdamW optimizer:
$\min_{v_{*}}\mathcal{L}=\begin{cases}\mathcal{L}_{b},&\text{if the input has benign intent.}\\ \mathcal{L}_{m},&\text{if the input has malicious intent.}\end{cases}$ (3)
5 Evaluation

### v3

@@ Abstract
Recent text-to-image (T2I) models have exhibited remarkable performance in generating high-quality images from text descriptions. However, these models are vulnerable to misuse, particularly generating not-safe-for-work (NSFW) content, such as sexually explicit, violent, political, and disturbing images, raising serious ethical concerns. In this work, we present PromptGuard, a novel content moderation technique that draws inspiration from the system prompt mechanism in large language models (LLMs) for safety alignment. Unlike LLMs, T2I models lack a direct interface for enforcing behavioral guidelines. Our key idea is to optimize a safety soft prompt that functions as an implicit system prompt within the T2I model’s textual embedding space. This universal soft prompt ($P^{*}$) directly moderates NSFW inputs, enabling safe yet realistic image generation without altering the inference efficiency or requiring proxy models. We further enhance its reliability and helpfulness through a divide-and-conquer strategy, which optimizes category-specific soft prompts and combines them into holistic safety guidance. Extensive experiments across five datasets demonstrate that PromptGuard effectively mitigates NSFW content generation while preserving high-quality benign outputs. PromptGuard achieves 3.8 times faster than prior content moderation methods, surpassing eight state-of-the-art defenses with an optimal unsafe ratio down to 5.84%.
I Introduction
Text-to-image (T2I) models, like Stable Diffusion [1], enable realistic image generation from text prompts. However, their misuse for generating not-safe-for-work (NSFW) content (e.g., sexual and violent images) raises significant ethical concerns [2, 3, 4], including the spread of harmful content like AI-generated child sexual abuse material [5] and politically manipulative imagery [6]. Effective defense mechanisms for T2I services are urgently needed.
Current NSFW safeguards fall into two categories: model alignment and content moderation. Model alignment (e.g., fine-tuning) directly modifies the T2I model to remove NSFW capabilities [7, 8, 9, 10, 11, 12], but can degrade performance on benign inputs [10, 13]. Content moderation uses external models to filter unsafe textual inputs [14] or visual outputs [15], or employs prompt modification using LLMs [16] to promote safer generation. While avoiding unintended removal of benign concepts, these methods add computational overhead. An efficient and robust content moderation framework remains a critical need.
@@ I Introduction
In this paper, we present PromptGuard, a novel T2I moderation technique that optimizes a soft prompt that works as a system prompt for safety to neutralize malicious contents in input prompts in an input-agnostic manner without affecting benign image generation quality and performance. As shown in Figure 1, our basic idea draws inspiration from the “system prompt” mechanism in LLMs, which has exhibited remarkable effectiveness in aligning output content with safe and ethical guidelines [17, 18] and our approach seeks to apply similar guidance in T2I settings.
However, designing PromptGuard is challenging from two perspectives: First, T2I models, unlike LLMs, lack a direct mechanism for implementing system prompts. They treat all textual input as user-generated content, requiring a novel approach to emulate the system-prompt mechanism within the T2I context. Second, the diverse nature of NSFW content—including categories such as violence, sexual explicitness, and political extremism—makes it difficult to design a single, universal safeguard.
To address the first challenge, we introduce a safety pseudo-word, optimized within the continuous embedding space of the T2I model’s text encoder. This soft prompt effectively steers both benign and NSFW prompts (e.g., “A painting of a woman, nude, sexy") away from regions associated with unsafe content. Moreover, we employ SDEdit[19] to transform unsafe images into safer counterparts, allowing PromptGuard to learn how to generate realistic, safe images from potentially harmful inputs. This approach contrasts with existing moderation methods[14, 15, 9] that often block or blur undesirable outputs. For the second challenge, we categorize NSFW content into four types: sexual, violent, political, and disturbing [20, 21]. Rather than attempting to create a single universal soft prompt, we adopt a divide-and-conquer strategy, optimizing separate soft prompts for each category and then combining them. This approach improves the reliability and robustness of the moderation system. To ensure PromptGuard’s efficacy without negatively affecting benign image generation, we apply a contrastive learning-based method that balances strong NSFW suppression with the preservation of image quality.
@@
The extensive experiments compared PromptGuard with eight state-of-the-art defense techniques on five benchmark datasets. Our evaluation validates six key aspects of PromptGuard: (1) Effectiveness: Achieved the lowest unsafe ratio (5.84%) in a natural language setting, outperforming all baselines. (2) Universality: Ranked in the top two across all four NSFW categories. (3) Adversarial Robustness: Outperformed all baselines in NSFW removal under three adversarial attacks. (4) Efficiency: 3.8x faster than previous moderation methods without extra computational cost. (5) Helpfulness: Provides realistic, safe content instead of merely blocking or blurring NSFW outputs (Figure 4). (6) Scalability: Demonstrates flexibility in adapting to new NSFW categories. We also discuss limitations, future work, and have open-sourced our code on website [^1] Project Website: https://prompt-guard.github.io/ to foster further research in AI ethics.
@@
  • New Technique: We introduce the application of the system prompt concept to T2I models, using soft prompt optimization to achieve effective and lightweight content moderation.
  • New Findings: Our comprehensive experiments across diverse datasets demonstrate PromptGuard’s effectiveness, universality, adversarial robustness, efficiency, helpfulness, and scalability.
II Related Work
II-A Content Moderation
@@
To ensure the safe use of T2I models, existing methods implement safety measures for both input and output. Latent Guard [22] filters input text by classifying embeddings, allowing only safe prompts to pass through. In contrast, Stable Diffusion V1.4’s default safety filter [15] detects and blocks any NSFW images at the output stage by blacking them out. POSI [16] fine-tunes a language model to rewrite unsafe prompts into safe alternatives before passing them to the diffusion model. Some methods focus on enhancing safety during the generation process itself. Safe Latent Diffusion [23] adjusts the diffusion process to steer the text-conditioned guidance vector away from unsafe areas in the embedding space. However, these approaches often require additional models or modifications, which add to computational cost. In contrast, PromptGuard introduces a soft prompt that efficiently directs the model towards safe outputs without relying on external models or process changes.
II-B Model Alignment
@@
Another line of work directly fine-tunes models to enhance safety, rather than relying solely on additional guardrails. ESD [7] fine-tunes the diffusion model to direct the generative process away from undesired concepts, while UCE [8] modifies the text projection matrices to erase specific concepts from the model. Additionally, SafeGen [9] optimizes the self-attention layers to eliminate unsafe concepts in a text-agnostic manner. However, these methods require either model retraining or parameter fine-tuning, which introduces significant computational costs. In PromptGuard, we propose a soft prompt approach that removes unsafe concepts effectively without modifying model parameters, ensuring lightweight safety alignment.
III Background
III-A Text-to-Image (T2I) Generation
@@
The success of denoising diffusion models, such as DDPM [24], has advanced text-to-image (T2I) models like Stable Diffusion (SD) and Latent Diffusion [25]. These models rely on text encoders that transform text prompts into latent embeddings, guiding the image generation process. The text is tokenized and mapped into a high-dimensional embedding space, which influences the image synthesis through cross-attention during diffusion. For instance, SD uses the CLIP text encoder, which improves upon the BERT encoder used in Latent Diffusion [26], benefiting from a larger training set (LAION-5B [27]) for more effective embeddings. The encoder’s intermediate layers play a crucial role in progressively building complex concepts throughout the diffusion process. Recent studies, like the Diffusion Lens [28], show that early layers capture basic objects, while deeper layers establish relationships between elements.
III-B System Prompt
A system prompt is a predefined instruction given to large language models (LLMs) to guide their behavior, tone, and responses, ensuring safety and mitigating risks such as bias or harmful outputs [29, 30]. By embedding ethical guidelines, system prompts prevent misleading responses without modifying model parameters [31]. They are lightweight and effective, requiring minimal computational overhead compared to complex model fine-tuning. Although widely studied in LLMs, system prompts have not been explored in text-to-image (T2I) models, where textual descriptions guide visual content generation. Unlike LLMs, T2I models face unique challenges in prompt engineering for visual outputs. While user prompts influence image generation, system prompts for ethical constraints and output refinement have not been fully explored. In this work, we integrate system prompt mechanisms into T2I models for NSFW content moderation using a soft prompt approach (see IV).
IV PromptGuard
IV-A Overview
Figure 2: Diagram of PromptGuard. The training data preparation consists of two types of data: (1) malicious prompts paired with images, including both the original malicious image and its edited, safer version, and (2) benign prompts paired with corresponding images. The individual soft prompt embedding training involves appending a trainable soft token embedding to the end of the original prompt token embeddings. Focusing on one unsafe category at a time, we train only the parameters of the soft token embedding using the loss function $L_{m}$ or $L_{b}$, depending on whether the input is benign or malicious. During inference, we concatenate all the trained embeddings and append them to the end of the user input, functioning as a soft system prompt.
@@ IV-A Overview
In this section, we introduce the design of PromptGuard, which aims to optimize a soft prompt suffix $P^{*}$ that is appended to user inputs for NSFW content moderation. This soft prompt has two primary objectives: (1) mitigating harmful semantics while preserving safe content in malicious prompts and (2) ensuring fidelity in benign image generation. Directly identifying an effective prompt suffix at the token level is challenging due to the discrete nature of text space. To overcome this, we optimize the soft prompt in the token embedding space, leveraging techniques from prompt tuning [32, 33] and prompt-driven safety mechanisms in LLMs [31], operating within a continuous domain.
To address the first objective, we employ contrastive learning, constructing training pairs where harmful content serves as negative data and its moderated counterpart as positive data. To address the second objective, adversarial training which incorporates benign data into the training dataset ensures that benign prompts remain unaffected, preserving the quality of benign image generation.
Since a single embedding vector may struggle to moderate diverse NSFW categories, we adopt a divide-and-conquer approach. Unsafe content is categorized into four types—sexual, violent, political, and disturbing—following prior classification frameworks [20, 21]. Each category is assigned a dedicated safe token embedding, which is optimized individually and later concatenated into a unified soft prompt. This composite embedding is consistently appended to all user inputs during inference, ensuring efficient, parameter-free content moderation. Figure 2 illustrates our training and inference pipeline.
IV-B Training Data Preparation
@@
To construct the training dataset for each specific unsafe category, we collect malicious prompts from three sources: the prompt dataset provided by UD [20], prompts created by a NSFW data scraper [34], and prompts generated by GPT-4o mini [35, 36]. Combining these sources ensures a diverse and comprehensive dataset. For benign prompts, we use GPT-4o [37] to generate examples across 6 categories: animals, food, human beings, landscapes, transport vehicles, home scenes, which are selected based on the COCO dataset [38], a widely used benchmark for diverse, real-world image generation. Since our goal is to find a soft prompt that can guide safe visual generation, we utilize the T2I model being safeguarded to create images based on these collected prompts.
@@
As outlined in  IV-A, we then construct safer versions of each malicious image by altering only the unsafe visual elements while leaving safe parts intact. To achieve these fine-grained modifications, we leverage SDEdit [19] to generate safer images based on the original malicious ones. This process requires specific prompts to guide the image-to-image generation. For sexually explicit content, we apply the simple prompt “Fully dressed person, safe for work.” For the other three unsafe categories, we employ GPT-4o to transform each malicious image’s prompt into a safe version, which then guides the image-to-image generation. This approach enables us to create a high-quality image pair dataset, as illustrated in Figure 3.
@@
IV-C Individual Soft Prompt Embedding Training
@@
Before adjusting $v_{*}$, the SD model’s encoder in the VAE module first transforms the image $x^{\text{ben}}$ or the image pair $x^{\text{org}},x^{\text{tgt}}$] into clean latent representations $z_{0}^{\text{ben}}$ or $z_{0}^{\text{org}},z_{0}^{\text{tgt}}$]. Then, the DDPM noise scheduler [24] iteratively injects noise $\epsilon_{t}^{\text{ben}}$ or $\epsilon_{t}^{\text{org}},\epsilon_{t}^{\text{tgt}}$] into the clean latent representations, resulting in noisy latent representations $z_{t}^{\text{ben}}$ or $z_{t}^{\text{org}},z_{t}^{\text{tgt}}$]. The denoising U-Net U takes both the noisy latent representation $z_{t}$, which contains visual information, and the hidden state embeddings $c$, which contain textual information, to predict the noise $\epsilon_{\text{U}}(z_{t},t,c)$ for the next $t$ steps. We aim for the model to correctly predict the noise added to the original latent representation, $\epsilon_{0}^{\text{ben}}$, given the condition $c_{b}$. Simultaneously, we want the model’s prediction, conditioned on $c_{m}$, to closely match $\epsilon_{t}^{\text{tgt}}$ while being far from $\epsilon_{t}^{\text{org}}$. This ensures that the model’s prediction is aligned with the noise added to the safer version of the image while becoming less accurate in predicting the noise for the original unsafe image. To achieve these two objectives, we design two separate loss functions: $\mathcal{L}_{b}$ (benign preservation) and $\mathcal{L}_{m}$ (malicious moderation). (1) For each benign input data:
@@
(2) For each malicious input data:
@@
Minimizing $L_{b}$ helps ensure that the prompt with our appended $P_{*}$ preserves the ability to correctly generate benign images. On the other hand, minimizing $\mathcal{L}_{m}$ encourages $P_{*}$ to guide the predicted noise to stay far from the original unsafe vision while becoming closer to the safe vision representations. The hyperparameter $\lambda$ controls the balance between these two objectives. Increasing $\lambda$ forces $P_{*}$ to focus more on keeping the model away from unsafe vision representations, reducing its ability to recover unsafe images from noise and encourage safe version generations. The overall optimization framework could be formalized using $\underset{v_{*}}{\min}\,\mathcal{L}$ as follows:
[…]
IV-D Inference
Once the individual safe embeddings for different NSFW categories (e.g., sexual, violent, political, disturbing) have been trained, they are concatenated into a unified composite soft prompt. This combined soft prompt is then appended to the end of every user input during inference, functioning as an implicit system prompt for the T2I model. Unlike traditional moderation techniques that rely on separate filtering models or prompt rewriting, this approach directly integrates safety guidance within the model’s textual embedding space, ensuring continuous, lightweight, and inference-efficient moderation.
V Experiments